OpenStudy (anonymous):

solve the proportion- 5z/3=105/6

OpenStudy (anonymous):

\[\frac{ 5z}{ 3 }=\frac{ 105}{ 6 }\]

OpenStudy (anonymous):

\[(-5z/3)\times6=(105/6)\times6\]

OpenStudy (anonymous):

-10z=105

OpenStudy (anonymous):

-10z/10=105/-10

OpenStudy (anonymous):

z=-1.5

OpenStudy (anonymous):

\[z=\frac{21}{2}=10.5 \]

OpenStudy (anonymous):

why is the answer negative?

OpenStudy (anonymous):

@mustry

OpenStudy (anonymous):

yap negative is answer

OpenStudy (anonymous):

coz question is -5z/3=105/6

OpenStudy (anonymous):

but why? cuz in the beginning it's a positive 5z/3...so you gotta switch it to a negative or something?

OpenStudy (anonymous):

ok i gt u question is 5z/3=105/6?

OpenStudy (anonymous):

oh i see what happened :p thnks man

OpenStudy (anonymous):

ok ll be 1.5

OpenStudy (anonymous):

got it xD